It is known that 35% of the mice inoculated with a serum are protected from certain disease. If 7 mice are inoculated, find the probability that at most 1 mouse survive the disease.
Answer #1

X ~ Binomial (n,p)

Where n = 7 , p = 0.35

P(X) = nCx px (1-p)n-x

P( X <= 1) = P (X = 0) + P (X = 1)

= 7C0 0.350 0.657 + 7C1 0.351 0.656

= 0.2338